# Physics-informed neural networks with unknown measurement noise

### v4

@@ 1 Introduction
Figure 1: The PINN-EBM approach: the inputs $t$ are passed through the PINN to obtain the PINN predictions $\hat{x}$. The differential operator $\mathcal{F}(\lambda)$ is then applied to $\hat{x}$ to obtain the PDE residuals, which are subsequently utilized to calculate the PDE loss $\mathcal{L}_{\rm PDE}$. At the same time, the residuals between the noisy measurements $y$ and $\hat{x}$ are formed, serving as noise estimates. The EBM is then trained on these estimates in order to learn the noise PDF, which can in turn be utilized to compute the likelihood of the measurements serving as data loss $\mathcal{L}_{\rm data}^{\rm EBM}$. Finally, the two loss terms are combined to form the total loss $\mathcal{L}_{\rm tot}$. Both PINN and EBM, as well as the PDE parameters $\lambda$, can be trained by backpropagating $\mathcal{L}_{\rm tot}$. to not mess up formatting
In this work, we consider the inverse problem for the case of measurements contaminated with non-Gaussian noise of unknown form. The least-squares loss, which is commonly employed as data loss in PINNs, is known to perform poorly in this case (Constable, 1988; Akkaya and Tiku, 2008). We give a way of mitigating this issue by suitably modeling the noise distribution. A high-level illustration of our method is given in Fig. 1: we employ an energy-based model (EBM) to learn the noise probability density function (PDF) jointly with the PINN. This PDF is then utilized to estimate the likelihood of the measurements under our model, which in turn serves as data loss.
Non-Gaussian measurement noise of unknown form can appear in a variety of applications (Johnson and Rao, 1991). Measurements of geomagnetic fields may exhibit asymmetric, long-tailed noise (Constable, 1988), data in astrophysics or seismology may be contaminated with impulsive noise (Weng and Barner, 2005), and not accounted for systematic errors in the measurement procedure may give rise to bias in the noise (Barlow, 2002). For example, the noise distribution considered in Fig. 1 can be interpreted as resulting from the case where measurements of multiple sensors have been merged into one dataset. One of the sensors produces biased measurements, giving rise to the second peak away from zero. Then the PINN-EBM allows for solving the PDE correctly together with the simultaneous detection of previously unrecognized systematic errors in the data.
@@ 2 Related work
Some research has been done on PINNs in case of noisy measurements, although typically only Gaussian noise is considered. In Yang et al. (2021), the framework of Bayesian neural networks (Goan and Fookes, 2020) is combined with the PINN framework, in order to obtain uncertainty estimates when training PINNs on noisy data. Bajaj et al. (2021) introduce the GP-smoothed PINN, where a Gaussian process (GP) (Rasmussen and Williams, 2006) is utilized to ameliorate noisy initial value data and make the PINN training more robust in this situation. In Chen et al. (2021), the PINN-SR method is introduced which can be employed to determine governing equations from scarce and noisy data. In contrast to these works, we give a way of taking into account unknown, non-Gaussian noise in the PINN and provide a training procedure for this case.
@@ 3 Background
In this section we give a brief introduction to the two distinct methods which we will combine in our work: physics-informed neural networks and energy-based models.
@@ 3.2 Energy-based models (EBMs)
In our case, the evaluation of the partition function to high accuracy remains tractable by utilizing numerical integration, since the noise distribution is one-dimensional (note that this does not constitute a restriction on the dimensionality of the PDE).
@@ 4.1 PINNs in case of non-zero mean noise
For noise with non-zero mean, however, the problem of mismatch between the two loss terms is not simply an effect of finite data, which could be treated via an adequate regularization procedure, but has deeper roots. Considering again the limit of infinite data, we have $\underset{N_{d}\rightarrow\infty}{\lim}\mathcal{L}_{\rm data}=\mathbb{E}[(\hat{x}-x-\epsilon)^{2}]=\mathbb{E}[(\hat{x}-x-\mu_{\epsilon})^{2}]+\text{Var}_{\epsilon}$, and the minimizer of the data loss becomes $\hat{x}=x+\mu_{\epsilon}$, where $\mu_{\epsilon}$ denotes the mean of the noise distribution. For the PDE loss, on the other hand, the minimizer would remain $\hat{x}=x$ (assuming that the correct parameters $\lambda$ of $\mathcal{F}(\lambda)$ are known).
@@ 4.4 Training PINN and EBM jointly
Initializing the EBM only later during the training process is advantageous for two reasons: firstly, the least-squares loss is still a sensible first guess, so initializing the EBM on the residuals stemming from the pretrained network will let it start out with a reasonable form of the likelihood instead of a random one. Secondly, it will give us a better idea on the range in which the residuals between PINN prediction and data lie, allowing for a better normalization of the inputs to the EBM and hence more efficient training.
@@
When evaluating the performance of our models, the following metrics are considered: the absolute error in the learned values of the PDE parameters ($|\Delta\lambda|$), the root-mean-square error (RMSE) on the validation data, the negative log-likelihood (NLL) of the validation data according to the models, and the square values of the PDE residuals ($f^{2}$) on the training data. In practice, the NLL is the most relevant performance metric: contrary to the RMSE and $|\Delta\lambda|$, it can also be calculated when the true solution is not known. While $f^{2}$ can also be calculated without knowledge of the true solution, it only measures how well the learned PDE is fulfilled and not necessarily the correct one.
Implementation details and further results are discussed in Appendix A and B: the method is tested on an additional PDE, and the model performance is investigated as a function of various dataset properties and training parameters.
@@ 5.1 Toy problem
3G $100|\Delta\lambda|$ 1.22$\pm$1.12 3.96$\pm$3.62 10.22$\pm$1.47
RMSE 1.89$\pm$1.44 6.06$\pm$6.82 8.27$\pm$1.85
NLL 3.48$\pm$0.97 7.85$\pm$9.00 8.88$\pm$2.75
$100f^{2}$ 0.51$\pm$0.30 30.85$\pm$11.51 48.83$\pm$17.07
u $100|\Delta\lambda|$ 1.27$\pm$0.33 3.25$\pm$2.45 12.29$\pm$0.66
RMSE 1.90$\pm$0.83 3.98$\pm$3.67 9.09$\pm$1.09
NLL 4.75$\pm$1.14 6.49$\pm$6.77 19.06$\pm$3.82
$100f^{2}$ 0.72$\pm$0.29 16.25$\pm$4.75 37.94$\pm$12.56
G $100|\Delta\lambda|$ 2.45$\pm$1.78 2.02$\pm$1.34 1.08$\pm$0.91
RMSE 3.37$\pm$2.06 3.23$\pm$2.65 2.83$\pm$2.47
NLL 5.13$\pm$2.47 5.05$\pm$3.55 4.15$\pm$2.48
$100f^{2}$ 0.42$\pm$0.32 10.82$\pm$5.97 11.72$\pm$6.58
3G0 $100|\Delta\lambda|$ 1.17$\pm$1.03 5.28$\pm$3.23 2.21$\pm$2.03
RMSE 2.22$\pm$1.40 7.90$\pm$6.89 5.00$\pm$5.06
NLL 3.73$\pm$1.29 9.76$\pm$11.15 5.44$\pm$3.99
$100f^{2}$ 0.61$\pm$0.41 32.20$\pm$15.05 34.02$\pm$15.53
@@
In Table 1, performance metrics of the different models as obtained for different noise forms are given. The results have been averaged over ten runs with different realizations of the measurement noise.
We start by discussing the results for Gaussian mixture noise (3G). It is evident that PINN performs poorly on all metrics. While PINN-off learns the correct parameter $\lambda$ better on average, the variance of its predictions is very high, thereby limiting its use. The PINN-EBM estimate of the parameter is significantly less variable than that of PINN-off, and is very close to the correct value of $\lambda$. In accordance with the quality of the parameter estimates, PINN-EBM clearly outperforms both PINN and PINN-off when considering the RMSE and logL; it achieves the lowest RMSE as well as the highest logL on the validation data.
@@
In Fig. 3, the impact of various training parameters on the model performance is shown, when considering 3G noise (left plot). It is apparent that the weighting parameter $\omega$ can significantly impact the model performance. From the two plots on the right, we observe that the performances of the three models converge as either the number of training points or the noise strength is reduced; for higher values, the impact of non-Gaussian noise is larger and PINN-EBM outperforms. More in-depth discussions of these results can be found in Appendix B.
@@ 5.2 Navier-Stokes equations
[Uncaptioned image] Figure 4: The NLL on the validation data as a function of the parameter $\omega$, weighting the PDE loss. The averages of 5 runs are depicted, plus-or-minus one standard deviation.
noise PINN-EBM PINN-off PINN 3G $100|\Delta\lambda_{1}|$ 1.19$\pm$0.67 2.92$\pm$0.47 23.10$\pm$0.11 $100|\Delta\lambda_{2}|$ 0.04$\pm$0.03 0.09$\pm$0.05 0.08$\pm$0.06 RMSE 0.06$\pm$0.01 0.11$\pm$0.01 0.20$\pm$0.02 NLL -0.03$\pm$0.08 0.15$\pm$0.07 0.40$\pm$0.30 $100f^{2}$ 0.04$\pm$0.00 0.10$\pm$0.01 0.18$\pm$0.01
Table 2: Results for the Navier Stokes equations (9a) in case of Gaussian mixture (3G) noise. The entries in the table are averages obtained over 5 runs plus-or-minus one standard deviation. Bold font highlights best performance.
The results are given in Table 2. It is apparent that PINN-EBM learns $\lambda_{1}$ better than PINN and PINN-off whereas the difference is small for $\lambda_{2}$. PINN-EBM also performs better in terms of RMSE and NLL. Here, we have chosen $\omega=50$ for PINN-EBM since the NLL on the validation data decreased when the PDE loss was weighted more strongly (see Fig. 4). For PINN and PINN-off, on the other hand, increasing the parameter did not lead to notable changes in the NLL and we kept $\omega=1$.
@@ 6 Conclusions and future work
In this paper, we demonstrated that the standard PINN fails in case of non-zero mean measurement noise and we proposed the PINN-EBM to resolve this problem; utilizing an EBM to learn the noise distribution allows the PINN to produce good results also in case of non-zero mean and non-Gaussian noise. Using several examples, ranging from a simple toy problem to the complex Navier-Stokes equations, we demonstrated the capabilities of our method and showed that it outperforms the standard PINN by a significant margin in case of non-Gaussian noise. In addition to determining the correct PDE solution, the PINN-EBM also allows for the identification of the true noise distribution, which may result in novel insights into the measurement procedure.
In the future, it would be interesting to investigate combining the PINN-EBM with other improvements to the PINN framework, such as adaptive weighting schemes (McClenny and Braga-Neto, 2020; Xiang et al., 2022), scheduling approaches (Krishnapriyan et al., 2021; Wang et al., 2022) or additional loss terms (Yu et al., 2022). In principle, our method could also be applied to the problem considered in Bajaj et al. (2021) (see also Section 2): where the standard GP presumably would fail here as well in case of non-Gaussian noise, a representation for the initial conditions could potentially also be learned via PINN-EBM. Finally, it would be interesting to apply the method to suitable real-world datasets.
@@ A.1 Additional experiment: the Bessel equation
To test the framework on a more complicated differential equation than the exponential equation (8), we employ the Bessel equation, a second-order differential equation (Niedziela, 2008; Bowman, 2012),
@@
where we pick $\nu=1$ and where we have introduced the parameter $\lambda=0.7$, which is to be estimated by the PINN. Results for one run of this experiment are shown in Fig. 5. When comparing the results obtained here (see Table 3) to those from the previous section, we note the following: PINN-EBM outperforms both PINN and PINN-off on all metrics except $f^{2}$, and PINN-off performs better than PINN. For this example, the variance of the results for both PINN and PINN-off is smaller than in the previous experiment. One reason for this can be found in the $t^{2}$ term in (10), which can lead to large values in the PDE loss in (6), thereby putting more weight on PDE fulfillment than on matching the data to a Gaussian.
@@
[Uncaptioned image] Figure 5: The Bessel equation (10) with Gaussian mixture noise (3G). Results for one individual run, where the blue dots represent training data and the red dots validation data.
noise PINN-EBM PINN-off PINN 3G $100|\Delta\lambda|$ 0.27$\pm$0.22 1.81$\pm$0.93 3.11$\pm$1.58 RMSE 0.00$\pm$0.00 0.04$\pm$0.02 0.06$\pm$0.03 NLL -0.94$\pm$0.13 -0.63$\pm$0.08 -0.19$\pm$0.32 $100f^{2}$ 0.63$\pm$0.14 0.39$\pm$0.40 0.19$\pm$0.05

### v5

@@ 1 Introduction
Figure 1: The PINN-EBM approach: the inputs $t$ are passed through the PINN to obtain the PINN predictions $\hat{x}$. The differential operator $\mathcal{F}(\lambda)$ is then applied to $\hat{x}$ to obtain the PDE residuals, which are subsequently utilized to calculate the PDE loss $\mathcal{L}_{\rm PDE}$. At the same time, the residuals between the noisy measurements $y$ and $\hat{x}$ are formed, serving as noise estimates. The EBM is trained on these estimates in order to learn the noise PDF, which can in turn be utilized to compute the likelihood of the measurements serving as data loss $\mathcal{L}_{\rm data}^{\rm EBM}$. Finally, the two loss terms are combined to form the total loss $\mathcal{L}_{\rm tot}$. Both PINN and EBM, as well as the PDE parameters $\lambda$, can be trained by backpropagating $\mathcal{L}_{\rm tot}$. to not mess up formatting
In this work, we consider the inverse problem for the case of measurements contaminated with homogeneous non-Gaussian noise of unknown form. The least-squares loss, which is commonly employed as data loss in PINNs, is known to perform poorly in this case (Constable, 1988; Akkaya and Tiku, 2008). We give a way of mitigating this issue by suitably modeling the noise distribution. A high-level illustration of our method is given in Fig. 1: we employ an energy-based model (EBM) to learn the noise probability density function (PDF) jointly with the PINN. This PDF is utilized to estimate the likelihood of the measurements under our model, which in turn serves as data loss.
Non-Gaussian measurement noise of unknown form can appear in a variety of applications (Johnson and Rao, 1991). Measurements of geomagnetic fields may exhibit asymmetric, long-tailed noise (Constable, 1988), data in astrophysics or seismology may be contaminated with impulsive noise (Weng and Barner, 2005), and not accounted for systematic errors in the measurement procedure may give rise to bias in the noise (Barlow, 2002). For example, the noise distribution considered in Fig. 1 can be interpreted as resulting from the case where measurements from multiple sensors have been merged into one dataset. One of the sensors produces biased measurements, giving rise to the second peak away from zero. Then the PINN-EBM allows for solving the PDE correctly together with the simultaneous detection of previously unrecognized systematic errors in the data.
@@ 2 Related Work
Prior research has been done on PINNs in case of noisy measurements, although typically only Gaussian noise is considered. In Yang et al. (2021), the framework of Bayesian neural networks (Goan and Fookes, 2020) is combined with the PINN framework, in order to obtain uncertainty estimates when training PINNs on noisy data. Psaros et al. (2023) discuss different methods of uncertainty quantification and Zou et al. (2023) consider the case where both inputs and outputs are noisy. Bajaj et al. (2023) introduce the GP-smoothed PINN, where a Gaussian process (GP) (Rasmussen and Williams, 2006) is utilized to ameliorate noisy initial value data and make the PINN training more robust in this situation. In Chen et al. (2021), the PINN-SR method is introduced which can be employed to determine governing equations from scarce and noisy data. In contrast to these works, we give a way of taking into account unknown, non-Gaussian noise in the PINN and provide a training procedure for this case.
@@ 3 Background
In this section we give a brief introduction to the two distinct methods which we combine in our work: physics-informed neural networks and energy-based models.
@@ 3.2 Energy-based Models (EBMs)
In our case, the evaluation of the partition function to high accuracy remains tractable by utilizing numerical integration, since the noise distribution is one-dimensional. Higher-dimensional outputs are still possible, as long as the noise is uncorrelated. Also, note that this does not constitute a restriction on the dimensionality of the inputs $t$ of the PDE.
@@ 4.1 PINNs in Case of Non-zero Mean Noise
For noise with non-zero mean, however, the problem of mismatch between the two loss terms is not simply an effect of finite data, which could be treated via an adequate regularization procedure, but has deeper roots. Considering again the limit of infinite data, we have $\underset{N_{d}\rightarrow\infty}{\lim}\mathcal{L}_{\rm data}=\mathbb{E}[(\hat{x}-x-\epsilon)^{2}]=\mathbb{E}[(\hat{x}-x-\mu_{\epsilon})^{2}]+\text{Var}_{\epsilon}$, and the minimizer of the data loss becomes $\hat{x}=x+\mu_{\epsilon}$, where $\mu_{\epsilon}$ denotes the mean of the noise distribution. For the PDE loss, on the other hand, the minimizer remains $\hat{x}=x$ (assuming that the correct parameters $\lambda$ of $\mathcal{F}(\lambda)$ are known).
@@ 4.4 Training PINN and EBM Jointly
Initializing the EBM only later during the training process is advantageous for two reasons: firstly, the least-squares loss is still a sensible first guess, so initializing the EBM on the residuals stemming from the pretrained network will let it start out with a reasonable form of the likelihood instead of a random one. Secondly, it will give us a better idea about the range in which the residuals between the PINN prediction and the data lie, allowing for a better normalization of the inputs to the EBM and hence more efficient training.
@@
When evaluating the performance of our models, the following metrics are considered: the absolute error in the learned values of the PDE parameters ($|\Delta\lambda|$), the root-mean-square error (RMSE) between $\hat{x}$ and $x$ in the range of the validation data, the negative log-likelihood (NLL) of the validation data according to the models, and the square values of the PDE residuals ($f^{2}$) in the range of the training data. In practice, the NLL is the most relevant performance metric: contrary to the RMSE and $|\Delta\lambda|$, it can also be calculated when the true solution is not known. While $f^{2}$ can also be calculated without knowledge of the true solution, it only measures how well the learned PDE is fulfilled and not necessarily the correct one.
Additional results are discussed in Appendices A and B: the method is tested on another PDE, and the model performance is investigated as a function of various dataset properties and training parameters.
@@ 5.1 Toy Problem
3G $100|\Delta\lambda|$ 1.2$\pm$1.1 4.0$\pm$3.6 10.2$\pm$1.5
RMSE 1.9$\pm$1.4 6.1$\pm$6.8 8.3$\pm$1.9
NLL 3.5$\pm$1.0 7.9$\pm$9.0 8.9$\pm$2.8
$100f^{2}$ 0.5$\pm$0.3 30.9$\pm$11.5 48.8$\pm$17.1
u $100|\Delta\lambda|$ 1.3$\pm$0.3 3.3$\pm$2.5 12.3$\pm$0.7
RMSE 1.9$\pm$0.8 4.0$\pm$3.7 9.1$\pm$1.1
NLL 4.8$\pm$1.1 6.5$\pm$6.8 19.1$\pm$3.8
$100f^{2}$ 0.7$\pm$0.3 16.3$\pm$4.8 37.9$\pm$12.6
G $100|\Delta\lambda|$ 2.5$\pm$1.8 2.0$\pm$1.3 1.1$\pm$0.9
RMSE 3.4$\pm$2.1 3.2$\pm$2.7 2.8$\pm$2.5
NLL 5.1$\pm$2.5 5.1$\pm$3.6 4.2$\pm$2.5
$100f^{2}$ 0.4$\pm$0.3 10.8$\pm$6.0 11.7$\pm$6.6
3G0 $100|\Delta\lambda|$ 1.2$\pm$1.0 5.3$\pm$3.2 2.2$\pm$2.3
RMSE 2.2$\pm$1.4 7.9$\pm$6.9 5.0$\pm$5.1
NLL 3.7$\pm$1.3 9.8$\pm$11.2 5.4$\pm$4.0
$100f^{2}$ 0.6$\pm$0.4 32.2$\pm$15.1 34.0$\pm$15.5
@@
In Table 1, performance metrics of the different models as obtained for different noise forms and with $\omega=1$ are given. The results have been averaged over ten runs with different realizations of the measurement noise.
We start by discussing the results for Gaussian mixture noise (3G). It is evident that PINN performs poorly on all metrics. While PINN-off learns the correct parameter $\lambda$ better on average, the variance of its predictions is very high, thereby limiting its use. The PINN-EBM estimate of the parameter is significantly less variable than that of PINN-off, and is very close to the correct value of $\lambda$. In accordance with the quality of the parameter estimates, PINN-EBM clearly outperforms both PINN and PINN-off when considering the RMSE and NLL; it achieves the lowest RMSE as well as the lowest NLL on the validation data.
@@
In Fig. 3, the impact of various training parameters and dataset properties on the model performance is shown, when considering 3G noise (left plot). It is apparent that the weighting parameter $\omega$ can significantly impact the model performance. From the two plots on the right, we observe that the performances of the three models converge as either the number of training points or the noise strength is reduced; for higher values, the impact of non-Gaussian noise is larger and PINN-EBM outperforms. More in-depth discussions of these results can be found in Appendix B.
@@ 5.2 Navier-Stokes Equations
The results are given in Table 2. It is apparent that PINN-EBM learns $\lambda_{1}$ better than PINN and PINN-off whereas the difference is smaller for $\lambda_{2}$. PINN-EBM also performs better in terms of RMSE and NLL. Here, we have chosen $\omega=50$ for PINN-EBM since the NLL on the validation data decreased when the PDE loss was weighted more strongly (see Fig. 4). For PINN and PINN-off, increasing the parameter did not lead to notable changes in the NLL and we kept $\omega=1$.
5.3 Implementation Details
Figure 4: The NLL on the validation data as a function of the parameter $\omega$, weighting the PDE loss. The averages of 5 runs are depicted, plus-or-minus one standard deviation.
noise PINN-EBM PINN-off PINN
3G $100|\Delta\lambda_{1}|$ 1.19$\pm$0.67 2.92$\pm$0.47 23.10$\pm$0.11
$100|\Delta\lambda_{2}|$ 0.04$\pm$0.03 0.09$\pm$0.05 0.08$\pm$0.06
RMSE 0.06$\pm$0.01 0.11$\pm$0.01 0.20$\pm$0.02
NLL -0.03$\pm$0.08 0.15$\pm$0.07 0.40$\pm$0.30
$100f^{2}$ 0.04$\pm$0.00 0.10$\pm$0.01 0.18$\pm$0.01
Table 2: Results for the Navier Stokes equations (9) in case of Gaussian mixture (3G) noise. The entries in the table are averages obtained over 5 runs plus-or-minus one standard deviation. Bold font highlights best performance.
For both PINN and EBM, fully-connected neural networks with tanh activation function were used. Both inputs and outputs of the networks were normalized to their expected ranges. For the PINN, 4 layers of width 40 were chosen for the exponential equation; in case of the Navier-Stokes equations, 8 layers of width 20 were used instead. For the EBM, 3 layers of width 5 were used and a dropout layer with factor $0.5$ was inserted before the last layer.
The PINN was trained for $40.000$ iterations in case of the exponential equation (Table 1) and $100.000$ iterations for the Navier-Stokes equations (Table 2). When evaluating the impact of training parameters (Figs. 3, 4), $50.000$ iterations were used for both of them. The EBM was initialized for $N_{\rm EBM}=2000$ iterations, after $i_{\rm EBM}=4000$ and $i_{\rm EBM}=10000$ iterations, respectively, for the exponential equation and the Navier-Stokes equations. Convergence was checked by confirming that the values of the PDF tended toward zero on both ends of the distribution; if the check failed, the initialization was repeated with an increased number of iterations until it passed.
The datasets employed in Section 5.1 contained 200 training points, 50 validation points and 2000 collocation points. For the Navier-Stokes example in Section 5.2, 4000 training points, 1000 validation points and 4000 collocation points were used. The collocation points were generated on a uniform grid. The Adam optimizer with a learning rate of $0.002$ was used for both PINN and EBM. In case of the Navier-Stokes equations, the learning rate was reduced by a factor of $0.3$ after 80 0008000080\,00080 000 iterations. The batch size for data points was 200 (i.e. full batches were used for the exponential equation) and the batch size for collocation points was 100. Training the PINN-EBM typically took 30-50% longer than the standard PINN.
@@ 6 Conclusions and Future Work
In this paper, we demonstrated that the standard PINN fails in case of non-zero mean measurement noise and we proposed the PINN-EBM to resolve this problem; utilizing an EBM to learn the homogeneous noise distribution allows the PINN to produce good results also in case of non-zero mean and non-Gaussian noise. Using several examples, ranging from a simple toy problem to the complex Navier-Stokes equations, we demonstrated the capabilities of our method and showed that it outperforms the standard PINN by a significant margin in case of non-Gaussian noise. In addition to determining the correct PDE solution, the PINN-EBM also allows for the identification of the true noise distribution, which may result in novel insights into the measurement procedure.
In the future, it would be interesting to investigate combining the PINN-EBM with other improvements to the PINN framework, such as adaptive weighting schemes (McClenny and Braga-Neto, 2020; Xiang et al., 2022), adaptive collocation point placement (Wu et al., 2023), scheduling approaches (Krishnapriyan et al., 2021; Wang et al., 2024) or additional loss terms (Yu et al., 2022). Extending the approach to certain kinds of heteroscedastic noise and investigating the applicability of the approach to other frameworks such as neural ODEs (Chen et al., 2018) constitute promising avenues for future research. Finally, it would be interesting to apply the method to suitable real-world datasets.
@@ A.1 Additional Experiment: the Bessel Equation
To test the framework on a more complicated differential equation than the exponential equation (8), we employ the Bessel equation, a second-order differential equation (Niedziela, 2008; Bowman, 2012):
@@
where we pick $\nu=1$ and where we have introduced the parameter $\lambda=0.7$, which is to be estimated by the PINN. Results for one run of this experiment are shown in Fig. 5 and results for multiple runs are given in Table 3. When comparing the results obtained here to those from Section 5.1, we note the following: PINN-EBM outperforms both PINN and PINN-off on all metrics except $f^{2}$, and PINN-off performs better than PINN. For this example, the variance of the results for both PINN and PINN-off is smaller than for the exponential equation. One reason for this can be found in the $t^{2}$ term in (10), which can lead to large values in the PDE loss in (6), thereby putting more weight on PDE fulfillment than on matching the data to a Gaussian.
@@
Figure 5: The Bessel equation (10) with Gaussian mixture noise (3G). Results for one individual run, where the blue dots represent training data and the red dots validation data.
noise PINN-EBM PINN-off PINN
3G $100|\Delta\lambda|$ 0.27$\pm$0.22 1.81$\pm$0.93 3.11$\pm$1.58
RMSE 0.00$\pm$0.00 0.04$\pm$0.02 0.06$\pm$0.03
NLL -0.94$\pm$0.13 -0.63$\pm$0.08 -0.19$\pm$0.32
$100f^{2}$ 0.63$\pm$0.14 0.39$\pm$0.40 0.19$\pm$0.05